# Denis Roio Dyne.org &Rebecca Selvaggini

### v1

@@ 5.2 Future development
Key-value claims
The focus of the initial version of the SD-BLS algorithm is on claims consisting of fixed and predetermined values (e.g.: "is above 18" or "has M.Sc. in applied cryptography"), whereby a credential could contain multiple claims. This initial version doesn’t include the possibility to process claims containing a predetermined part (the key) and a variable one (the value). We’re expecting to be able support key-value based claims in the next version of the paper.
@@
EUDI-ARF dictates that the holder’s secret key generation and signatures must occur inside a trusted platform module (TPM). For mobile devices, this limits the secret keys and signatures to those offered, via proprietary APIs by the mobile OS, namely RSA (multiple flavours) and ECDSA on the secp256r1 curve. Currently the TPMs APIs supported by Android and iOS do not support BLS 12381 key generation or signature. Client-side signatures in EUDI-ARF are mostly used in the authentication process, specifically in the proof of possession required by the OpenID4VCI[19] issuance flow, but not in the verification. Therefore, we can investigate the possibility to use SD-BLS to implement a partially retro-compatible superset of EUDI-ARF, by maintaining the current issuance and verification protocols and using an extended SD-JWT format. Also, because of the unlinkable feature of credentials and revocations in SD-BLS, we can explore useful integrations with the European Blockchain Services Infrastructure (EBSI [20]).
@@
The SD-BLS math is fully compatible with ETH2.0 and can be computed inside an Ethereum VM. A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [22]) that want to authenticate peers using on a set of universally recognizable credentials instead of a key based proof of possession.

### v2

@@ 5.2 Future development
EUDI-ARF dictates that the holder’s secret key generation and signatures must occur inside a trusted platform module (TPM). For mobile devices, this limits the secret keys and signatures to those offered, via proprietary APIs by the mobile OS, namely RSA (multiple flavours) and ECDSA on the secp256r1 curve. Currently the TPMs APIs supported by Android and iOS do not support BLS 12381 key generation or signature.
Client-side signatures in EUDI-ARF are mostly used in the authentication process, specifically in the proof of possession required by the OpenID4VCI[17] issuance flow, but not in the verification.
Therefore, we can investigate the possibility to use SD-BLS to implement a partially retro-compatible superset of EUDI-ARF, by maintaining the current issuance and verification protocols and using an extended SD-JWT format. Also we can explore useful integrations with the European Blockchain Services Infrastructure (EBSI [18]).
Signroom and DIDroom
In SD-BLS both credential issuers and revocation issuers are in charge of various interactive administrative operations, while the dealing of revocation shares can be easily automated.
We plan to integrate SD-BLS in the free and open source software “Signroom”, an application we developed in the context of the NGI ASSURE grant, and “DIDroom” the dashboard connected to our did:dyne W3C DID domain.
@@
The SD-BLS math is fully compatible with ETH2.0 and can be computed inside an Ethereum VM. A verifier implemented in solidity can be a building block for more advanced Distributed Autonomous Organizations (DAO [20]) that want to authenticate peers using the selective disclosure of verifiable credentials instead of a key based proof of possession.
6 Acknowledgements
This work has been funded by the EU in the framework of the NGI TRUSTCHAIN project, grant No 101093274. We thank Puria Nafisi Azizi and Matteo Cristino for their infatiguable work on Zenroom, Luca Di Domenico for his help on the PVSS implementation, Giuseppe De Marco for sharing findings from his pioneering journey, Simone Onofri for his insights on threat models and Gabriele Bellini for his helpful security review. We extend our thanks the anonymous reviewers for their valuable advice and the De Cifris association for facilitating connections over a large network of unique professionals.